OpenStudy (anonymous):

Are these expressions polynomial functions? Why or why not? 1) 3x^8-6x^2+(pi)(times)x 2) 3x^8-6ix^2+(pi)(times)x

OpenStudy (anonymous):

they are not functions because they don't show any relations between x and y

OpenStudy (anonymous):

So what would be an example of a polynomial function?

OpenStudy (anonymous):

both of them are polynomial functions.

OpenStudy (anonymous):

Thank you! I thought so..... I wasn't sure about #2 because it had two variables

OpenStudy (anonymous):

@chenna, why r they polynomial functions

OpenStudy (anonymous):

they r polynomial expressions, not functions

OpenStudy (anonymous):

It says,"Cross out the expressions that are not polynomial functions."

OpenStudy (anonymous):

and what exactly does that mean?

OpenStudy (anonymous):

what do you mean by ' a function' ? define a function.

OpenStudy (anonymous):

They both are polynomial functions. function is a mapping, i.e. which maps values from one set to another set. for ex. y = 2*x, z=log(x) etc. sometimes we write y=2*x and sometimes f(x) =2*x. both are same. there is no difference. And the functions given above are polynomials. Hence they are polynomial functions.

OpenStudy (anonymous):

So can a polynomial function have more than one variable?

OpenStudy (anonymous):

what other variable do you see, apart from 'x' in the above equations ?

OpenStudy (anonymous):

i

OpenStudy (anonymous):

yes, a function is a mapping. But in that "function" I see no other set. I only see x as a variable

OpenStudy (anonymous):

a function is say y=x. It is a relation that says that for all values of x, y is equal to that x. It maps x to y

OpenStudy (anonymous):

So then what's the "i" for in the second option?

OpenStudy (anonymous):

it can be the complex root, the square root of -1

OpenStudy (anonymous):

'i' is not a variable. its just another constant. To explain you better, consider the polynomial \[x^{2}+1\]. = \[x ^{2}-i ^{2}\]. First function maps x in to space of real values and 2nd function maps x into complex space.